OpenStudy (clamin):

Please help!! Fill in the blanks with two consecutive integers to complete the following inequality. ____ < √181 < ____

OpenStudy (astrophysics):

\[\sqrt{181} \approx 13.45\], now think of two numbers that are close to that, which you can make an inequality out of.

OpenStudy (anonymous):

Since \(\sqrt x\) is an increasing function, you have \[\sqrt{169}<\sqrt{181}<\sqrt{196}\]and we didn't just pull those numbers out of thin air - their square roots are consecutive natural numbers.

OpenStudy (astrophysics):

Yes, I should've said it should be represented as such :P, and I thought it was obvious it was consecutive numbers as it is stated. Just giving them an idea of what it was, and figured they knew what a consecutive integer was. :)

OpenStudy (anonymous):

No harm done, it's hard to say whether or not OP has access to a calculator for this question or not.

OpenStudy (astrophysics):

You're right, thanks Siths!